OpenStudy (anonymous):

Let f(x)=0 for x not 1/n and let f(1/n)=1. Show that integral of f from 0 to1 is 0

OpenStudy (anonymous):

f(x)=1 ,for \[x \neq \frac{ 1 }{ n }\] and \[f (\frac{ 1 }{n })=1\].

OpenStudy (anonymous):

show \[\int\limits_{0}^{1}f(x)=0\]